OpenStudy (anonymous):

What is the inverse of y = 5x + 20?

OpenStudy (anonymous):

switch the y and x values, x = 5y + 20 Solve for y, x-20 = 5y y=(x-20)/5 Therefore the inverse function is y=(x-20)/5

OpenStudy (anonymous):

thank you !

OpenStudy (anonymous):

yw